### v1

Abstract
After pre-training on extensive image-text pairs, Contrastive Language-Image Pre-training (CLIP) demonstrates promising performance on a wide variety of benchmarks. However, a substantial volume of non-paired data, such as multimodal interleaved documents, remains underutilized for vision-language representation learning. To fully leverage these unpaired documents, we initially establish a Real-World Data Extraction pipeline to extract high-quality images and texts. Then we design a hierarchical retrieval method to efficiently associate each image with multiple semantically relevant realistic texts. To further enhance fine-grained visual information, we propose an image semantic augmented generation module for synthetic text production. Furthermore, we employ a semantic balance sampling strategy to improve dataset diversity, enabling better learning of long-tail concepts. Based on these innovations, we construct RealSyn, a dataset combining realistic and synthetic texts, available in three scales: 15M, 30M, and 100M. Extensive experiments demonstrate that RealSyn effectively advances vision-language representation learning and exhibits strong scalability. Models pre-trained on RealSyn achieve state-of-the-art performance on multiple downstream tasks. To facilitate future research, the RealSyn dataset and pre-trained model weights are released at https://github.com/deepglint/RealSyn. footnotetext: * Equal contribution. footnotetext: $\dagger$ Corresponding author.
RealSyn: An Effective and Scalable Multimodal Interleaved Document Transformation Paradigm
[…]
1 Introduction
The rapid proliferation of mobile networks and social platforms has led to exponential growth in large-scale data, offering a robust foundation for vision-language representation learning Guo et al. (2019); Baltrušaitis et al. (2018); Guo et al. (2024). By predicting the correspondence between images and description texts, Contrastive Language-Image Pre-training (CLIP) Radford et al. (2021) pretrains separate uni-modal encoders and achieves excellent performance across a range of downstream tasks Tian et al. (2020); He et al. (2020); Chen et al. (2020); Chopra et al. (2005); Shao et al. (2024); Sun et al. (2024); Luo et al. (2024). Notably, the representational capacity of such models improves significantly with the scale of the training dataset Li et al. (2024b). However, the availability of paired data in the real world is limited.
Figure 1: Multimodal interleaved documents are unsuitable for vision-language representation learning. We construct distinct image-text pairs from such documents via retrieval and generation.
In recent years, the release of large-scale image-text pair datasets Schuhmann et al. (2021, 2022); Thomee et al. (2016); Byeon et al. (2022); Gadre et al. (2024) has significantly advanced vision-language representation learning. Due to download failures and non-English captions in the YFCC Thomee et al. (2016) dataset, DeCLIP Li et al. (2022) reprocesses a new version of the YFCC15M dataset. LAION400M Schuhmann et al. (2021) provides 400 million image-text pairs collected from the web and filtered using CLIP similarity. COYO700M Byeon et al. (2022) offers 747 million high-quality pairs curated through advanced filtering strategies, including CLIP similarity, watermark detection, and aesthetic scoring. Despite these advancements, a substantial amount of non-paired data, particularly interleaved image-text documents Zhu et al. (2024); Laurençon et al. (2024); Li et al. (2025) containing multiple images and text paragraphs without explicit correspondence, remains incompatible with conventional vision-language representation learning methods.
Recent studies Fan et al. (2024); Yu et al. (2024); Zheng et al. (2024a); Lai et al. (2024) aim to enhance the quality of open-source image-text datasets using existing models. For instance, LaCLIP Fan et al. (2024) leverages large language models (LLMs) to rewrite raw captions for better alignment with images. CapsFusion Yu et al. (2024) fine-tunes an LLM using a ChatGPT-generated instruction dataset to mitigate hallucination issues. However, the diversity and distribution of synthetic data generated by these methods are inherently constrained by the limitations of the underlying generative models.
Figure 2: The Real-World Data Extraction pipeline to extract high-quality images and texts from interleaved image-text documents.
In this paper, we explore two fundamental questions: 1) How to utilize multimodal interleaved documents for vision-language representation learning. 2) How to effectively leverage both realistic and synthetic texts to enhance representation performance. To this end, as shown in Fig. 1, we first establish a Real-World Data Extraction pipeline to extract high-quality images and texts. Then we design a hierarchical retrieval method to efficiently associate each image with multiple semantically relevant texts. To enhance fine-grained image understanding, we propose a visual semantic augmented generation module for synthetic text production. Furthermore, we employ a semantic balance sampling strategy to improve dataset diversity, enabling better learning of long-tail concepts. Based on these innovations, we construct the RealSyn dataset, which integrates both realistic and synthetic texts with three sizes: 15M, 30M, and 100M. Comprehensive experimental results show that RealSyn is effective for vision-language representation learning and exhibits excellent scalability. The main contributions of this paper are summarized as follows:
  • We propose an effective and scalable multimodal interleaved document transformation paradigm for vision-language representation learning.
  • We release RealSyn, a large-scale semantic balanced dataset that integrates both realistic and synthetic texts and is available in three sizes: 15M, 30M, and 100M.
  • We conduct extensive experiments and demonstrate the effectiveness and scalability of our proposed RealSyn dataset.
2 Related Work
Large-Scale Pre-training Dataset. In recent years, several large-scale image-text datasets Chen et al. (2015); Clark and Gardner (2017); Goyal et al. (2017); Byeon et al. (2022); Li et al. (2024a) collected from the Internet have been released. The YFCC100M Thomee et al. (2016) dataset provides a comprehensive overview of the evolution of photo and video documentation and sharing from the inception of Flickr in 2004 until early 2014. Due to download failures and non-English captions, DeCLIP Li et al. (2022) reprocesses a new version of the YFCC15M dataset. Additionally, the LAION400M Schuhmann et al. (2021) dataset contains 400 million image-text pairs collected from Common Crawl and widely used in vision-language pre-training. Recent advancements have also introduced several large-scale interleaved image-text document datasets Li et al. (2025); Zhu et al. (2024); Laurençon et al. (2024). The OBELICS Laurençon et al. (2024) dataset uses a comprehensive filtering strategy and includes 141 million web pages, 353 million associated images, and 115 billion text tokens extracted from Common Crawl. However, due to data format constraints and training inefficiencies, interleaved image-text documents are currently unsuitable for vision-language representation learning.
Vision Language Pre-training. As a pioneering work in visual language pre-training, CLIP has attracted extensive attention due to its powerful zero-shot recognition and exceptional transfer learning performance Wang and Kang (2024); Tang et al. (2024); Shao et al. (2024); Martin et al. (2024). Inspired by CLIP, numerous visual-language pre-training works have been published in recent years Mu et al. (2022); Li et al. (2022); Yang et al. (2023). SLIP Mu et al. (2022) enhances performance by combining self-supervised learning with CLIP pre-training. DeCLIP Li et al. (2022) improves pre-training efficiency by integrating multi-view supervision across modalities and nearest-neighbor supervision from similar pairs. To mitigate the influence of noisy data, ALIP Yang et al. (2023) introduces a gating mechanism to dynamically allocate weights to samples. However, all these methods rely on large-scale image-text pairs crawled from the Internet.
Synthetic Captions. Recent works Yang et al. (2023); Yu et al. (2024); Chen et al. (2024) indicate that image-text pairs obtained from websites contain intrinsic noise, which directly impacts the effectiveness of vision-language pre-training. To enhance data diversity, LaCLIP Fan et al. (2024) uses the in-context learning capability of large language models to rewrite text descriptions associated with each image. CapsFusion Yu et al. (2024) employs large language models to refine information from web-based image-text pairs and synthetic captions, improving the quality of multimodal pre-training data. However, the quality and diversity of synthetic data are inherently limited by the capabilities of the generative model.
3 RealSyn Dataset
3.1 Real-World Data Extraction
To transform interleaved image-text documents for vision-language representation learning, we establish a Real-World Data Extraction pipeline (Fig. 2) to extract high-quality images and texts. This pipeline consists of three steps: Data Extraction, Image Filtration, and Sentence Filtration.
Data Extraction. We employ 118M interleaved image-text documents from the OBELICS Laurençon et al. (2024) as the primary data source. All images are extracted and stored in a dedicated image database, while sentences are segmented using the Natural Language Toolkit (NLTK) Bird and Loper (2004) and stored in a separate sentence database. This process yields 336M images and 2.13B sentences from the interleaved documents.
Image Filtration. After extracting 336M images, we apply a two-stage filtering process to ensure data quality and reduce redundancy. First, we discard images that meet any of the following criteria: 1) the shorter dimension is fewer than 100 pixels, or 2) the aspect ratio exceeds 3 or is below 1/3. This step removes 51M low-quality images. Next, following CLIP-CID Yang et al. (2025), we use the EVA02-CLIP E/14-plus model Sun et al. (2023) to extract image embeddings and apply the Union-Find algorithm Tarjan (1975) to eliminate perceptually and semantically redundant images. This step removes an additional 87M images, resulting in a refined dataset of 198M high-quality images.
Sentence Filtration. After extracting 2.13B sentences from interleaved image-text documents, we conduct rigorous filtering based on quality, semantics, and redundancy. Initially, we eliminate sentences based on the following criteria: 1) presence of emojis or URLs; 2) sentences containing fewer than 3 or more than 81 words; and 3) following CAT Radenovic et al. (2023), we retain samples with at least C1 caption complexity and incorporating an action. This phase reduces the corpus size from 2.13B to 1.82B sentences. Then we apply semantic filtering to the remaining sentences, eliminating those with minimal information assessed through information entropy:
$\theta(x)=-\sum_{i=1}^{n}p(x_{i})\log p(x_{i}),$ (1)
where $n$ denotes the number of words in a sentence, $x_{i}$ represents the $i$-th words in the sentences $x$, and $p(x_{i})$ is the probability of the word $x_{i}$ in the entire corpus. Based on human cognition principles and empirical experience, we filter out sentences with a score below 0.3. To further refine the corpus by removing difficult or ambiguous sentences, we use GTP2-large Radford et al. (2019) to calculate the perplexity score $\mathcal{PPL}$ for each sentence:
$\mathcal{PPL}(x)=\exp\left\{-\frac{1}{t}\sum_{i=1}^{t}\log p_{\theta}(x_{i}\mid x_{<i})\right\},$ (2)
where $t$ represents the token number of the sentence, and $p_{\theta}(x_{i}\mid x_{<i})$ is the likelihood of the $i$-th token given the previous tokens. We retain sentences with perplexity scores between 30 and 200. The overall semantics filtering reduces the corpus to 1.16B sentences. In the final stage, similar to redundancy image filtering, we perform both perceptual and semantic deduplication of sentences. This process results in a refined corpus of 0.84B sentences that include extensive real-world knowledge.
Figure 3: The architecture of our proposed framework, which constructs distinct image-text pairs from real-world data extracted from interleaved documents via retrieval and generation.
3.2 Retrieval and Generation Framework
After extracting high-quality images and sentences from documents, we propose an efficient and scalable framework to retrieve multiple semantically relevant texts for each image and leverage large language models to integrate retrieved realistic text with fine-grained visual information and generate synthetic text. As shown in Fig. 3, the architecture of our framework primarily consists of three components: Text Semantic Clustering, Hierarchical Retrieval, and Image Semantic Augmented Generation.
Data Scale Dataset
Food101
CIFAR10
CIFAR100
Birdsnap
SUN397
Cars
Aircraft
DTD
Pets
Caltech
Flowers
STL10
EuroSAT
RESISC45
KITTI
Country
UCF101
Memes
SST2
ImageNet
Average
15M YFCC 67.2 90.4 70.8 47.7 66.7 23.8 29.7 62.4 65.7 80.1 90.0 94.7 94.9 79.4 75.4 18.4 70.8 48.6 56.2 56.7 64.5
LAION 71.0 93.3 78.1 41.0 66.3 76.9 43.0 71.2 74.5 87.6 88.2 93.6 95.3 82.9 72.2 13.5 75.4 55.7 57.3 59.3 69.8
RealSyn 77.1 94.5 78.7 43.4 71.4 64.7 42.7 71.3 79.9 90.0 88.2 96.4 96.2 87.2 72.4 16.7 79.9 55.7 57.7 64.0 71.4
30M LAION 76.1 94.5 80.0 47.4 70.3 82.3 45.9 74.7 80.3 89.8 89.5 95.6 95.5 84.5 72.6 15.2 76.6 56.2 60.0 64.3 72.6
RealSyn 81.2 95.4 81.8 48.4 74.5 73.4 45.2 74.2 84.1 91.3 90.6 97.2 96.5 89.2 74.5 19.0 82.6 55.0 56.2 68.5 73.9
100M LAION 80.2 95.7 82.5 51.3 73.4 85.3 46.1 75.6 83.2 91.1 92.0 96.9 95.2 85.9 68.4 17.4 80.0 57.3 61.4 68.3 74.4
RealSyn 84.2 96.3 83.5 54.0 76.2 77.4 47.6 75.6 86.3 92.1 91.7 97.7 96.8 90.6 73.1 21.1 83.7 57.3 58.9 71.6 75.8
Table 1: Linear probe on 20 downstream datasets. Pre-training ViT-B/32 on RealSyn achieves 1.3%-6.9% average performance improvement.
Data Scale Dataset
Food101
CIFAR10
CIFAR100
Birdsnap
SUN397
Cars
Aircraft
DTD
Pets
Caltech
Flowers
STL10
EuroSAT
RESISC45
KITTI
Country
UCF101
Memes
SST2
ImageNet
Average
15M YFCC 36.3 74.0 40.3 19.4 41.8 2.1 2.3 12.0 19.8 59.8 48.9 87.7 21.2 20.3 23.8 5.1 27.8 47.4 50.1 32.3 33.6
LAION 49.1 85.7 56.9 11.5 45.1 49.9 3.8 25.7 54.6 78.1 30.5 89.5 36.7 36.1 21.7 5.6 38.2 48.8 49.9 37.1 42.7
RealSyn 60.0 85.7 58.3 10.5 56.4 27.6 5.5 33.2 61.7 80.2 31.2 92.4 56.5 56.2 34.0 8.9 52.6 53.3 51.3 43.3 47.9
30M LAION 58.9 85.9 63.1 17.4 54.8 61.0 4.3 36.4 65.5 82.0 41.3 91.3 40.3 43.7 24.3 7.2 47.4 51.5 50.1 44.9 48.6
RealSyn 67.5 89.0 65.2 15.0 60.6 39.2 7.9 37.8 70.5 84.0 42.2 93.8 59.9 61.9 27.7 10.6 56.7 52.5 50.1 50.9 52.1
100M LAION 68.9 90.5 68.6 23.6 60.6 68.3 7.8 41.2 74.7 87.1 47.7 94.4 45.6 53.4 23.6 10.4 54.5 51.9 53.3 52.8 53.9
RealSyn 73.5 89.5 68.8 20.1 65.0 48.5 10.2 46.1 76.7 87.6 48.8 94.4 69.0 65.5 24.6 12.1 60.5 52.4 54.1 56.2 56.2
Table 2: Zero-shot transfer on 20 downstream datasets. Pre-training ViT-B/32 on RealSyn achieves 2.3%-14.3% average performance improvement.
Text Semantic Clustering. To efficiently retrieve multiple semantic relevant texts for each image, we initially encode all sentences using the EVA02-CLIP E/14-plus Sun et al. (2023) model. Inspired by Unicom An et al. (2023), we utilize the standard $K$-Means algorithm Ikotun et al. (2023) offline to partition the 0.84B texts into 2M clusters with the help of efficient feature quantization Johnson et al. (2019).
Hierarchical Retrieval. Given the prohibitive computational overhead of direct semantic text retrieval from 0.84B sentences (exceeding 10,000 hours on 8 × A100 GPUs), we devise a hierarchical retrieval method to optimize computational efficiency. We first perform inter-cluster retrieval to find the most relevant cluster center for each image. Then, we group images sharing the same cluster center and perform intra-cluster retrieval to obtain multiple semantically relevant sentences. This approach enables the retrieval of 198M images and 0.84B sentences within 40 hours using 8 $\times$ A100 GPUs.
Image Semantic Augmented Generation. Although the retrieved realistic texts achieve satisfactory performance, they exhibit limitations in capturing fine-grained visual semantics. To address this issue, we introduce the Image Semantic Augmented Generation module. This module initially employs the OFA model to generate a concise caption for each image. We then integrate the open-set image tagging model RAM++, which extracts object detection tags. Considering that RAM++ supports only 4,000 tags, we expand this set to 8,000 by incorporating an additional 4,000 tags derived from real-world sentences. Following CapsFusion Yu et al. (2024), we utilize ChatGPT4 Turbo to merge the retrieved realistic texts with concise captions and image tags to construct a 100K instruction dataset (the prompt we used is presented in the supplementary material). Subsequently, we fine-tune the LLaMA3-8B model Dubey et al. (2024) using the LLaMA Factory Zheng et al. (2024b) and deploy the vLLM Kwon et al. (2023) for large-scale inference. Ultimately, we convert data from 118M multimodal interleaved documents into 198M image-text pairs, where each image is associated with multiple retrieved realistic texts and synthetic texts.
3.3 Semantic Balance Sampling
To further improve the quality and diversity of our dataset, we implement semantic balancing sampling across the 198M image-text pairs. Specifically, we use EVA02-CLIP E/14-plus Sun et al. (2023) to encode and calculate the cosine similarity between imaged and synthetic texts. To reduce the impact of OCR-related or mismatched pairs during pre-training, we filter out 29.7M pairs with cosine similarities above 0.61 or below 0.51. Inspired by MetaCLIP Hu Xu (2023), we introduce an easy but efficient cluster-based semantic balance sampling strategy. We cluster the image embeddings from the remaining 168.3M pairs into 1M centers. To enhance the semantic diversity of our dataset, we randomly select 20, 35, and 180 samples from clusters exceeding these thresholds, while retaining all samples from smaller clusters. This approach culminates in the construction of the RealSyn15M, RealSyn30M, and RealSyn100M datasets.
4 Experiments and Results
4.1 Implementation Details
We initially collect 118M interleaved image-text documents from the OBELICS Laurençon et al. (2024) as our primary data source. We use OFAbase Wang et al. (2022) and RAM++large Huang et al. (2023) to generate brief captions and semantic tags. To validate the dataset performance, we pre-train standard CLIP supervised by the text randomly selected from the three retrieved realistic texts and one synthetic text. During pre-training, we adopt AdamW Loshchilov (2019) as the optimizer, with a learning rate of 1e-3 and a weight decay of 0.2. The parameters $\beta 1$ and $\beta 2$ are set to 0.9 and 0.98, respectively. The input image size is 224×224, and the input text sequence length is 77. The temperature parameter $\tau$ is initialized to 0.07. We train 32 epochs with 4096 batch sizes on 8 $\times$ A100 (80G) GPUs. Please refer to the supplementary material for more details.
To validate the effectiveness of the RealSyn dataset, we compare RealSyn with the previous datasets across various models and data scales. We compare RealSyn15M with the YFCC15M filtered by DeCLIP Li et al. (2022). Following ALIP Yang et al. (2023), we also compare with LAION15M, LAION30M, and LAION100M (subset randomly selected from LAION400M).
Text Retrieval Image Retrieval
Flickr30k MSCOCO Flickr30k MSCOCO
Data Scale Dataset R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10
15M YFCC 37.1 64.8 75.9 21.3 45.1 57.0 23.5 47.3 58.3 13.2 32.0 43.1
LAION 49.1 76.8 84.5 28.4 53.0 64.9 33.3 60.5 70.9 17.4 38.3 49.7
RealSyn 72.9 91.1 95.1 43.8 69.5 79.6 49.5 76.3 84.6 25.8 50.6 62.5
30M LAION 59.6 83.5 89.8 35.9 62.4 73.2 42.4 70.1 79.4 22.1 45.5 57.6
RealSyn 76.0 93.3 96.9 48.2 74.6 83.0 54.0 80.0 87.6 29.5 55.2 66.9
100M LAION 67.5 87.9 93.0 43.3 68.0 78.1 50.4 77.2 85.5 27.1 52.1 63.8
RealSyn 81.6 96.1 97.3 52.3 76.7 85.0 58.8 84.1 90.5 32.5 58.9 70.2
Table 3: Zero-shot image-text retrieval performance on Flickr30k and MSCOCO. Pre-training CLIP-B/32 on RealSyn dataset achieves a significant improvement on all metrics.
4.2 Main Results
Linear Probe. In Tab. 1, we present the linear probe performance of the ViT-B/32 model across 20 downstream datasets. When pretrained on the 15M scale, RealSyn15M surpasses YFCC15M in 16 out of 20 datasets, registering an average performance increase of 6.9%. Moreover, RealSyn15M outperforms LAION15M in 18 out of 20 datasets, with an average improvement of 1.6%. Upon scaling the dataset to 30M and 100M, RealSyn achieves average performance improvement of 1.3% and 1.4% over LAION, respectively. These results demonstrate the efficacy of the RealSyn dataset in vision-language representation learning.
Zero-shot Transfer. We evaluate the zero-shot transfer performance of the ViT-B/32 model across 20 classification benchmarks using the same prompt templates as SLIP Mu et al. (2022). As indicated in Tab. 2, RealSyn15M surpasses YFCC15M on 18 of 20 datasets, achieving an average performance improvement of 14.3%. In comparison to LAION15M, RealSyn15M excels on 18 of 20 datasets with an average improvement of 5.2%. Upon expanding the dataset sizes to 30M and 100M, RealSyn achieves average performance improvements of 3.5% and 2.3% compared to LAION, highlighting its efficiency and scalability.
It is important to note that RealSyn demonstrates a significant decrease in performance on certain datasets, such as Cars and Flowers. This reduction is primarily attributed to the unique data distribution of RealSyn, characterized by a scarcity of data for specific concepts, which hampers the model’s ability to effectively learn these concepts. For instance, as depicted in Fig. 4, samples related to cars constitute merely 0.9% of the dataset.
Data Scale Dataset IN-V2 IN-A IN-R ObjectNet IN-Sketch Average
15M YFCC 27.3 12.3 20.8 25.3 6.3 18.4
LAION 30.7 6.0 46.5 28.7 24.3 27.2
RealSyn 37.1 12.5 47.7 35.0 25.4 31.5
30M LAION 37.5 8.9 54.4 35.5 31.8 33.6
RealSyn 42.9 16.1 56.6 41.5 31.9 37.8
100M LAION 44.6 12.2 62.5 42.2 37.9 39.9
RealSyn 47.6 19.7 62.5 45.8 37.9 42.7
Table 4: Zero-shot robustness comparison. Pre-training CLIP-B/32 on RealSyn demonstrates superior robustness across all datasets.
Zero-shot Image-Text Retrieval. In Tab. 3, we present the zero-shot image-text retrieval performance of the ViT-B/32 model pre-trained on different scale of datasets. RealSyn achieves superior results across all evaluation metrics. Specifically, RealSyn15M improves Recall@1 by 35.8%&26% on Flickr30K Young et al. (2014) and by 22.5%&12.6% on MSCOCO Lin et al. (2014). RealSyn30M improves Recall@1 by 16.4%&11.6% on Flickr30K Young et al. (2014) and by 12.3%&7.4% on MSCOCO Lin et al. (2014). This significant enhancement in cross-modal retrieval performance indicates that the RealSyn dataset effectively improves vision-language representation learning by utilizing realistic and synthetic texts, resulting in robust representations and enhanced cross-modal alignment.
Figure 4: A T-SNE Van der Maaten and Hinton (2008) projection of LDA Blei et al. (2003) topic cluster from a randomly selected 1M samples from RealSyn.
Zero-shot Robustness. In Tab. 4, we present the zero-shot robustness performance. The results show that RealSyn significantly improves the robustness of vision-language pre-training models. Specifically, compared to LAION, RealSyn increases average performance by 4.3%, 4.2%, and 2.8% for datasets of 15M, 30M, and 100M, respectively. This notable improvement in performance primarily stems from the use of retrieved realistic texts that are not bound by the limitations of generative models, coupled with superior conceptual diversity compared to YFCC and LAION, thereby substantially enhancing model robustness.
5 Analysis
(a) Richness assessment comparison
(b) Diversity assessment comparison
Figure 5: The richness assessment and diversity assessment on different datasets. RealSyn-R1: the most relevant retrieved realistic text. RealSyn-S1: the semantic augmented synthetic text based on RealSyn-R1.
5.1 Statistics Analysis
Topic-based Assessment. Following MMC4 Zhu et al. (2024), we ran LDA Blei et al. (2003) on random sampling 1M image-realistic text pairs with 30 topics. Fig. 4 presents the proportions and examples for six topics: animal, food, airplane, flower, automotive, and landmark. Notably, the dataset contains minimal samples related to “flower” and “automotive” topics, representing merely 0.4% and 0.9% of the total, respectively. This paucity of examples hinders the model’s ability to sufficiently learn these concepts, thereby compromising its performance in the linear probe and zero-shot transfer evaluations on the Flowers and Cars datasets.
Richness Assessment. Fig. 5(a) presents image-text similarity and text token distribution of 15M samples from YFCC15, LAION, RealSyn-R1 (the most relevant retrieved realistic text), and RealSyn-S1 (the semantic augmented synthetic text based on RealSyn-R1). Compared to the datasets collected from the Internet, RealSyn exhibits robust similarity metrics, even after removing OCR data. Moreover, both the retrieved realistic texts and synthetic texts contain a larger quantity of words, which can provide a richer textual context that enhances vision-language representation learning.
Diversity Assessment. The RealSyn is constructed based on real-world interleaved image-text documents, which encompasses a wide array of diverse information. Following previous work Lai et al. (2025), we randomly select 0.2M samples to calculate the number of unique entities in the caption to assess the data diversity of different datasets. As depicted in Fig. 5(b), both the retrieved realistic texts and image semantic augmented synthetic texts exhibit a higher number of distinct entities. Such diversity enriches the dataset, facilitating the model’s acquisition of comprehensive knowledge and enhancing both performance and robustness.
Model Scaling. In Sec. 4.2, RealSyn exhibits superior performance across various data scales (More data scaling analysis is presented in the supplementary material). To further explore the model scaling capability, we present the downstream task performance of three models in Fig. 6. Notably, compared to LAION, RealSyn demonstrates steeper slopes in performance curves across linear probing, zero-shot transfer, and robustness, indicative of its superior model scaling capabilities.
Figure 6: Model scaling capability. We compare the models pre-trained on LAION30M and RealSyn30M.
Figure 7: Image captioning comparisons on COCO2017 and Flickr30k. B4, MT., RL. and Cd. represent the metric of BLEU, METEOR, ROUGE-L, and Cider.
5.2 Ablation Study
Image Captioning via MLLM. Following LLaVA-1.5 Liu et al. (2024), we initially align visual features to the text space using 558k data. Subsequently, we construct an image captioning dataset from LAION and RealSyn, which we then use for instruction tuning. Specifically, we split the realistic text and synthetic text for the same image into two independent samples, totaling 2M samples for training for one epoch. Concurrently, we randomly select 1M samples from LAION and train for two epochs. As illustrated in Fig. 7, RealSyn demonstrates a significant enhancement in performance compared to LAION across all evaluation metrics on both the COCO2017 Lin et al. (2014) and Flickr30k Young et al. (2014).
Extension to Pure Image. To further extend our method for pure images, we conduct experiments on ImageNet Deng et al. (2009). Initially, we retrieve semantically relevant realistic texts for each ImageNet image from our sentence database and generate image semantic augmented synthetic texts. Then, we pre-train ResNet50 He et al. (2016) supervised by the text randomly selected from the retrieved realistic texts and synthetic texts. Comparative analysis with SimCLR Chen et al. (2020) under identical conditions shows a linear probe average performance enhancement of 2.1% across 12 datasets using our constructed data. Detailed experimental results are provided in the supplementary material.
Model Dataset Linear probe Avg Transfer Avg Robustness Avg
CLIP-B/32 YFCC 64.5 33.6 18.4
LAION 69.8 42.7 27.2
RealSyn-Random 70.7 46.8 30.5
RealSyn-Balance 71.4 47.9 31.5
Table 5: Comparison of concept balance sampling and random sampling on the 15M dataset.
Ablation on Semantic Balance Sampling. To demonstrate the effectiveness of our proposed semantic balance sampling method, we compare it with random sampling. As indicated in Tab. 5, concept balance sampling yields performance improvements of 0.7%, 1.1%, and 1.0% in linear probe, zero-shot transfer, and robustness, respectively. Besides, we visualize the data distribution using different sampling methods by clustering 15M samples into 1M centers. As shown in Fig. 8, the distribution from semantic balanced sampling is smoother, facilitating the learning of long-tail concepts.
Figure 8: Clustering distribution of 15M data obtained from random sampling and semantic balanced sampling.
$T^{1}_{r}$ $T^{2}_{r}$ $T^{3}_{r}$ $T^{4}_{r}$ $T^{5}_{r}$ Linear probe Avg
70.3
71.0
71.2
70.9
70.6
$T^{1}_{s}$ $T^{2}_{s}$ $T^{3}_{s}$ $T^{4}_{s}$ $T^{5}_{s}$ Linear probe Avg
70.2
70.0
69.9
69.4
69.1
Table 6: Ablation experiment results using different types of text on the 15M dataset. $T^{k}_{r}$: the $k$-th retrieved semantic relevant realistic text. $T^{k}_{s}$: the image semantic augmented synthetic text for $T^{k}_{r}$.
Ablation on Realistic Texts and Synthetic Texts. We conduct ablation experiments on the impact of varying quantities of realistic and synthetic texts using the CLIP-B/32 model. As shown in Tab. 6, increasing the quantity of realistic text from 1 to 3 progressively improves model performance, attributable to enhanced text augmentation that provides extensive real-world knowledge. However, expanding from 3 to 5 slightly reduces performance due to information saturation and noise introduction. Meanwhile, increasing the number of synthetic text from 1 to 5 introduces more noise, gradually decreasing performance. Importantly, supervising image training solely with retrieved realistic texts significantly enhances performance, achieving a 71.2% accuracy compared to 69.8% with the LAION15M dataset. This highlights the crucial role of real-world knowledge in advancing vision-language representation learning.
Figure 9: Visualization of the raw interleaved document, the retrieved realistic text, and synthetic text. Image semantic-related information is highlighted in green.
Case Study. In Fig. 9, we present the visualization of retrieved realistic text and synthetic text obtained from an interleaved document using our proposed transformation paradigm. Both realistic and synthetic texts contain extensive descriptive information consistent with the image semantics, such as “lyell glacier”, “crags”, and “valley”. We provide more visualizations in the supplementary material.
6 Conclusion
This paper explores two open-ended questions: 1) How to utilize multimodal interleaved documents for vision-language representation learning. 2) How to effectively leverage both realistic and synthetic texts to enhance representation performance. To this end, we first establish a Real-World Data Extraction pipeline to extract high-quality images and texts. Then we design a hierarchical retrieval method to efficiently associate each image with multiple semantically relevant texts. To enhance fine-grained image understanding, we propose a visual semantic augmented generation module for synthetic text production. Furthermore, we employ a semantic balance sampling strategy to improve dataset diversity, enabling better learning of long-tail concepts. Based on these innovations, we present RealSyn, a dataset driven by both real and synthetic texts with three sizes: 15M, 30M, and 100M. Comprehensive experimental results show that RealSyn is efficient in vision-language representation learning and exhibits excellent scalability.
7 Limitations
To provide more fine-grained visual information, this paper utilizes vision expert models combined with a Large Language Model (LLM) to generate synthetic text. Given the inference cost and efficiency, the utilization of Modified Large Language Models (MLLM) for synthetic text generation is left for the community to explore. Furthermore, the transformation paradigm proposed in this paper is directly applicable to other multimodal document datasets, including MMC4 Zhu et al. (2024) and OmniCorpus Li et al. (2025). We hope our work provides insights into vision-language representation learning.
8 Ethics Statement
We abide by the ACL Code of Ethics. The data resources used in this study are publicly available.
Appendix A Detail Experiment Settings
A.1 Experiment Settings
In Tab. 7, we present the detailed settings used in training CLIP.
Hyperparameter Value
Initial temperature $0.07$
Weight decay $0.2$
Batch size 4096
Learning rate 0.001
Learning rate scheduler OneCycleLR
Pct start 0.1
Training epochs 32
GPU 8×8$\times$8 ×A100
Adam $\beta_{1}$ $0.9$
Adam $\beta_{2}$ $0.98$
Adam $\epsilon$ $10^{-6}$
Table 7: Hyperparameters used for CLIP pre-training.
A.2 Detail Instruction Prompt
The prompt we used for ChatGPT to construct the 100K instruction dataset is present in the following:
"Please merge the information from the given raw text and the synthetic caption with the help of the highly relevant detection tags. The raw caption offers detailed real-world information, yet it suffers from flaws in sentence structure and grammar. The synthetic caption exhibits impeccable sentence structure but often lacks in-depth real-world details and may contain false information. The highly relevant detection tags are provided to enrich the semantic information of the raw caption, while some are redundant and noisy. You are a great information integration and summary expert, you are also good at enriching semantic information. Ensure a well-structured sentence while retaining the detailed real-world information provided in the raw caption. Avoid simply concatenating the sentences and avoid adding external information to describe. Correctness and simplify sentences finally. Raw caption:<raw caption>, synthetic caption:<synthetic caption>, and highly relevant detection tags:<detection tags>".
Appendix B Detail External Results
Dataset Classes Train size Test size Evaluation metric
Food101 102 75,750 25,250 accuracy
CIFAR10 10 50,000 10,000 accuracy
CIFAR100 100 50,000 10,000 accuracy
Birdsnap 500 42,138 2,149 accuracy
SUN397 397 19,850 19,850 accuracy
Cars 196 8,144 8,041 accuracy
Aircraft 100 6,667 3,333 mean per class
DTD 47 3,760 1,880 accuracy
Pets 37 3,680 3,669 mean per class
Caltech101 101 3,000 5,677 mean per class
Flowers 102 2,040 6,149 mean per class
STL10 10 5,000 8,000 accuracy
EuroSAT 10 10,000 5,000 accuracy
RESISC45 45 3,150 25,200 accuracy
KITTI 4 6770 711 accuracy
Country211 211 42,200 21,100 accuracy
UCF101 101 9,537 1,794 accuracy
Memes 2 8,500 500 ROC AUC
SST2 2 7,792 1,821 accuracy
ImageNet 1000 1,281,167 50,000 accuracy
Table 8: List of linear probe datasets with the data distribution and evaluation metrics.
Model Dataset
Food101
CIFAR10
CIFAR100
Birdsnap
SUN397
Cars
Aircraft
DTD
Pets
Caltech
Flowers
STL10
EuroSAT
RESISC45
KITTI
Country
UCF101
Memes
SST2
ImageNet
Average
ViT-B/32 LAION 76.1 94.5 80.0 47.4 70.3 82.3 45.9 74.7 80.3 89.8 89.5 95.6 95.5 84.5 72.6 15.2 76.6 56.2 60.0 64.3 72.6
RealSyn 81.2 95.4 81.8 48.4 74.5 73.4 45.2 74.2 84.1 91.3 90.6 97.2 96.5 89.2 74.5 19.0 82.6 55.0 56.2 68.5 73.9
ViT-B/16 LAION 82.1 95.1 81.4 57.5 73.4 87.3 47.1 76.1 84.4 91.5 92.7 96.8 95.6 86.8 70.8 17.6 80.3 59.5 65.6 68.8 75.5
RealSyn 87.5 95.8 82.5 59.4 77.5 81.0 48.7 77.9 88.9 92.5 94.2 98.3 96.9 91.5 70.8 22.1 85.1 60.6 64.7 73.9 77.5
ViT-L/14 LAION 84.7 96.4 83.5 59.2 75.5 88.5 46.6 77.8 85.0 92.6 94.3 97.9 95.9 88.0 71.7 18.7 81.1 58.6 64.6 71.2 76.6
RealSyn 90.3 97.5 86.2 64.3 79.7 83.6 51.4 79.6 90.0 94.5 94.8 98.9 96.6 92.7 73.8 25.0 86.4 63.8 66.1 76.7 79.6
Table 9: Linear probe on 20 downstream datasets. Pre-training different scale CLIP models on RealSyn30M and LAION30M, achieves 1.3%-3.0% average performance improvement.
Model Dataset
Food101
CIFAR10
CIFAR100
Birdsnap
SUN397
Cars
Aircraft
DTD
Pets
Caltech
Flowers
STL10
EuroSAT
RESISC45
KITTI
Country
UCF101
Memes
SST2
ImageNet
Average
ViT-B/32 LAION 58.9 85.9 63.1 17.4 54.8 61.0 4.3 36.4 65.5 82.0 41.3 91.3 40.3 43.7 24.3 7.2 47.4 51.5 50.1 44.9 48.6
RealSyn 67.5 89.0 65.2 15.0 60.6 39.2 7.9 37.8 70.5 84.0 42.2 93.8 59.9 61.9 27.7 10.6 56.7 52.5 50.1 50.9 52.1
ViT-B/16 LAION 67.6 89.1 63.5 20.8 55.7 66.9 5.4 39.0 70.2 84.9 42.9 94.3 31.1 45.4 34.0 8.7 52.2 54.5 50.6 49.4 51.3
RealSyn 75.8 89.6 64.7 18.9 64.3 48.2 7.9 41.2 76.0 87.5 45.2 95.1 56.8 64.3 27.1 13.1 59.1 54.5 54.0 55.9 54.9
ViT-L/14 LAION 70.8 88.8 69.5 22.8 61.6 69.7 4.9 40.8 68.0 87.3 42.2 95.3 41.5 53.7 25.9 10.4 54.7 54.1 51.8 51.5 53.3
RealSyn 80.7 94.1 73.1 20.9 66.4 53.6 10.1 48.1 72.8 89.4 49.8 96.2 68.5 70.1 32.2 15.3 63.9 54.1 56.9 59.5 58.8
Table 10: Zero-shot transfer on 20 downstream datasets. Pre-training different scale CLIP models on RealSyn30M and LAION30M, achieves 3.5%-5.5% average performance improvement.
Model Dataset IN-V2 IN-A IN-R ObjectNet IN-Sketch Average
ViT-B/32 LAION 37.5 8.9 54.4 35.5 31.8 33.6
RealSyn 42.9 16.1 56.5 41.5 31.9 37.8
ViT-B/16 LAION 42.4 12.8 60.3 40.2 34.8 38.1
RealSyn 48.0 24.1 63.1 46.7 36.8 43.8
ViT-L/14 LAION 45.1 17.1 64.9 43.1 39.0 41.8
RealSyn 52.8 34.7 71.6 50.4 42.4 50.4
Table 11: Zero-shot robustness comparison. Pre-training different scale CLIP models on RealSyn30M and LAION30M, achieves 4.2%-8.6% average performance improvement.
B.1 Downstream Datasets
To comprehensively demonstrate the performance of CLIP trained on RealSyn, we compared the linear probe results of CLIP trained on RealSyn, YFCC Thomee et al. (2016), and LAION Schuhmann et al. (2021) across 20 datasets. These datasets include Food101 Bossard et al. (2014), CIFAR10 Krizhevsky et al. (2009), CIFAR100 Krizhevsky et al. (2009), Birdsnap Berg et al. (2014), SUN397 Xiao et al. (2010), Stanford Cars Krause et al. (2013), FGVC Aircraft Maji et al. (2013), DTD Cimpoi et al. (2014), Pets Parkhi et al. (2012), Caltech101 Fei-Fei et al. (2004), Flowers102 Nilsback and Zisserman (2008), SLT10 Coates et al. (2011), EuroSAT Helber et al. (2019), RESISC45 Cheng et al. (2017), KITTI Geiger et al. (2012), Country211 Radford et al. (2021), UCF101 Soomro et al. (2012), Hateful Memes Kiela et al. (2020), SST2 Radford et al. (2021), and ImageNet Deng et al. (2009). Details on each dataset and the corresponding evaluation metrics are provided in Tab. 8.
B.2 Detailed Model Scaling Results
Linear Probe. In Tab. 9, we present the detailed linear probe results of different scale CLIP model trained on the 30M dataset. The ViT-L/14 trained on RealSyn30M achieves an average performance improvement of 3.0% across 20 datasets compared to the model trained on the LAION30M.
Zero-shot Transfer. As shown in Tab. 10, ViT-L/14 trained on our proposed RealSyn30M outperforms LAION30M on 18 of 20 downstream datasets and achieves an average improvement of 5.5%.
Zero-shot Robustness. We present the detailed zero-shot robustness performance in Tab. 11, compared with LAION30M, the model trained on RealSyn30M boosts average robustness performance by 8.6% on ViT-L/14.
Figure 10: Data Scaling Analysis. Pre-training ViT-B/32 on RealSyn in different data scales.
B.3 Data Scaling Law
[…]
Model Method
Food101
CIFAR10
CIFAR100
Birdsnap
SUN397
Cars
Aircraft
VOC2007
DTD
Pets
Caltech101
Flowers
Average
ResNet50 SimCLR 68.4 90.6 71.6 37.4 58.8 50.3 50.3 80.5 74.5 83.6 90.3 91.2 70.6
Real&Syn Texts 72.5 89.1 69.0 57.1 63.6 51.4 48.1 85.5 69.7 90.5 88.1 88.8 72.7
Table 12: Linear probe on 12 downstream datasets. Pre-training CLIP (ResNet50) on image-text pairs achieves 2.1% performance improvement.
Dataset #Images #Avg Tokens /// Image #Avg Texts /// Image Text Type Source Type
CC12M 12 000 0001200000012\,000\,00012 000 000 -- 1 Realistic Website
YFCC15M 15 000 0001500000015\,000\,00015 000 000 16 1 Realistic Website
CapsFusion 120 000 000120000000120\,000\,000120 000 000 -- 1 Synthetic Image-Text Pair
LAION400M 400 000 000400000000400\,000\,000400 000 000 27 1 Realistic Website
RealSyn15M 15 239 4981523949815\,239\,49815 239 498 40 4 Realistic & Synthetic Interleaved Image-Text
RealSyn30M 30 328 8523032885230\,328\,85230 328 852 38 4 Realistic & Synthetic Interleaved Image-Text
RealSyn100M 100 862 786100862786100\,862\,786100 862 786 36 4 Realistic & Synthetic Interleaved Image-Text
Table 13: Current Dataset Comparison. Comparison with large-scale image-text pre-training datasets.
$T^{1}_{r}$ $T^{2}_{r}$ $T^{3}_{r}$ $T^{1}_{s}$ Linear probe Avg Transfer Avg Robustness Avg
70.3 42.4 25.7
71.2 46.8 30.7
70.2 39.7 24.0
71.4 47.9 31.5
Table 14: Ablation experiment results using different types of text on RealSyn15M. $T^{k}_{r}$: the $k$-th retrieved semantic relevant realistic text. $T^{k}_{s}$: the image semantic augmented synthetic text for $T^{k}_{r}$.
B.4 Ablation on Retrieved Realistic and Synthetic Texts
Tab. 14 shows the ablation experiment results on the text augmentation using different text types. After introducing the image semantic augmented synthetic text, due to its ability to supplement fine-grain visual semantic information, there are 0.2%, 1.1%, 0.8% performance enhancement on the linear probe, zero-shot transfer, and zero-shot robustness compared to solely utilizing the retrieved realistic texts.
B.5 Detailed Results on Pure Image
To further extend our method for pure images, we conduct experiments on ImageNet Deng et al. (2009). For each image, we retrieve three semantically relevant real-world sentences from our pre-constructed sentence database and generate a single semantically augmented synthetic caption based on the top retrieved text. Following SimCLR Chen et al. (2020), we utilize 4096 batch size and pre-train ResNet50 He et al. (2016) for 90 epochs supervised by the text randomly selected from the three retrieved realistic texts and one synthetic text.
As shown in Tab. 12, compared with SimCLR Chen et al. (2020) under the same conditions, the model trained on our constructed image-text pairs shows an average performance improvement of 2.1% across 12 downstream datasets. The results demonstrate that our method can effectively transform pure images into high-quality image-text pairs through retrieval and generation for vision-language pre-training.
Appendix C Further Analysis of RealSyn
C.1 Compare with Existing Datasets
In Tab. 13, we compare our proposed RealSyn dataset with existing widely used large-scale image-text pre-training datasets. Compared to previous datasets, our proposed RealSyn dataset provides four textual descriptions for each image, with an average token length of 36-40, significantly higher than LAION400M and YFCC15M. Furthermore, unlike previous datasets, the RealSyn dataset is sourced from real-world interleaved image-text documents and includes both realistic and synthetic texts, thereby expanding the scope for future research exploration.
C.2 Visualization of Examples
Figure 11: Visualization of image-text pairs in our proposed RealSyn dataset. $T^{k}_{r}$: the $k$-th retrieved realistic text. $T^{k}_{s}$: the image semantic augmented synthetic text for $T^{k}_{r}$. Image semantic-related information is highlighted in green.
Figure 12: Visualization of image-text pairs in our proposed RealSyn dataset. $T^{k}_{r}$: the $k$-th retrieved realistic text. $T^{k}_{s}$: the image semantic augmented synthetic text for $T^{k}_{r}$. Image semantic-related information is highlighted in green.
In Fig. 11 and Fig. 12, we visualize additional image-text pairs randomly selected from our proposed RealSyn dataset. $T^{k}_{r}$ is the $k$-th retrieved semantically relevant real-world sentence and $T^{k}_{s}$ is the semantic augmentic caption for $T^{k}_{r}$. We also highlighted the image semantic-related information in green and marked the image size below the image.